Z-Table Z Pr(Z) 0.38 65 0.50 69 0.67 75 0.84 80 1.04 85 1.28 90 92 1.41 1.56 94 1.65 95 1.75 96 98 2.06 2.33 99Chicagos Hard Rock Hotel distributes a mean of 800 bath towels per day to guests at the pool and in their rooms. This demand
Answer #1

Note: Please rate the answer (upvote) and do ask if any doubt is there.

Answer:

  • Reorder point = 3568 units
  • Safety stock = 368 units

Explanation:

  • Z value for 96% service level = 1.75
  • Demand during lead time = lead time * daily demand = 800*4 = 3200 units
  • Safety stock = Z * standard deviation of demand * (lead time)1/2 = 1.75*105*2 = 367.5 = 368 units (Ans B)
  • Reorder point = demand during lead time + Safety stock
  • Reorder point = 3200+ 367.5= 3567.5= 3568 units (Answer A)

Note: Please rate the answer (upvote) and do ask if any doubt is there.